OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −1/5(x − 2) y − 2 = −1/5(x + 3)

OpenStudy (anonymous):

Find the Slope First:

OpenStudy (anonymous):

\[m=\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

Then...Use Point Slope form

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

2 - 1 ------ -3 - 2 1 -- -5

OpenStudy (anonymous):

ok now i got it... y + 3 = −1/5(x − 2) (C)

OpenStudy (anonymous):

Nop.....this shuld be the answer y − 2 = −1/5(x + 3)

OpenStudy (anonymous):

Two point form: \[y-y1=m(x-x1)\]

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

(x-3) is with one set and (x+2) is with 2 sets

OpenStudy (anonymous):

Thanks with that i'm getting the others more easily.